Please someone could help me with this. Thanks in advance
If an={-3,2,7,12,17...}
Determine a138

Answer by rothauserc(4718)

this is an arithmetic sequence
an = a + d(n - 1), where d is the difference between terms and n is the nth term
In our example a = -3 and d = 5, therefore
a138 = -3 + 5(138-1) = 682